IF Y = 3X - 4 AND THE DOMAIN IS {-3,-1,4}, FIND THE RANGE
Answer by Ave(106)

domain is just another way of saying x, and range is y
plug in -3, -1, and 4 into the equation and solve for y for the range of each
----
for -3
y=3(-3)-4
y=0-4
y=-4
----
for -1
y=3(-1)-4
y= -3-4
y=-7
----
for 4
y=3(4)-4
y=12-4
y=8
range is {-4, -7, 8}